OpenStudy (anonymous):

Find the value of 53^3 using the identity (x+y)^3=x^3+3x^2y+3xy^2+y^2 Hint 53^3 =(50+3)^3; therefore, x=50, and y=3 I do not get this

OpenStudy (freckles):

replace x with 50 and y with 3

OpenStudy (freckles):

write x^3+3x^2y+3xy^2+y^2 but with 50 for x and 3 for y instead

OpenStudy (phi):

how much to you understand and what part is confusing ?

OpenStudy (anonymous):

(50+3^3= 1*50^3 + 3(50)^2(3) + 3(50)(3)^2 + 1*3^3= 125,000+22500+1350+27= 148,877

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@phi ??

OpenStudy (phi):

yes. and you can use a calculator and compute 53^3 directly. You should get the same answer

OpenStudy (anonymous):

okay thank you!!